Line D passes through the points (1,1) and (5,6). What is the slope of the line perpendicular to line D?
Answer by TimothyLamb(4379)

---
(1,1) (5,6)
---
to solve the linear equation, copy and paste this (the two given points):
1,1,5,6
into the "Two-Point form: x1 y1 x2 y2" input box here: https://sooeet.com/math/linear-equation-solver.php
---
x-intercept = (0.2, 0)
y-intercept = (0, -0.25)
slope = 1.25
---
answer:
perpendicular lines have negative reciprocal slopes, so the slope of the perpendicular line is:
perpendicular slope = -1/1.25 = -0.8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php